OpenStudy (anonymous):

Find the first six terms of the sequence. a1 = -1, an = 2 • an-1

OpenStudy (anonymous):

answers -2, -4, -8, -16, -32, -64 -1, -2, -4, -8, -16, -32 -1, -2, 0, 2, 4, 6 0, 2, -2, 0, 2, 4

OpenStudy (math&ing001):

a1 = -1 a2 = 2 . a1 = 2 . (-1) = -2 a3 = 2 . a2 = 2 . (-2) = -4 I'm sure you can do it from here.

OpenStudy (anonymous):

-2, -4, -8, -16, -32, -64

OpenStudy (jdoe0001):

\(\large { a_{\color{brown}{ n}}=2\cdot a_{{\color{brown}{ n}}-1}\qquad \begin{array}{lrllll} term& value \\\hline\\ a_1&-1\\ a_{\color{brown}{ 2}}&2\cdot a_{{\color{brown}{ 2}}-1}\to 2\cdot a_1\\ a_{\color{brown}{ 3}}&2\cdot a_{{\color{brown}{ 3}}-1}\to 2\cdot a_2\\ a_{\color{brown}{ 4}}&2\cdot a_{{\color{brown}{ 4}}-1}\to 2\cdot a_3\\ a_{\color{brown}{ 5}}&2\cdot a_{{\color{brown}{ 5}}-1}\to 2\cdot a_4\\ a_{\color{brown}{ 6}}&2\cdot a_{{\color{brown}{ 6}}-1}\to 2\cdot a_5\\ \end{array} }\)

OpenStudy (math&ing001):

Yes @amandaaaaaaaaaaa that's correct !